### v1

# Feature-Enhanced Neural Collaborative Reasoning with Dense Representation for Recommendation

Abstract
In recent years, various studies have focused on designing logical reasoning enhanced recommendation methods. However, previous studies either combine dense vector representation with relatively simple logical operators, or learn logical rule sets with one-hot encoding features. In this paper, we aim to simultaneously learn complex logical rules and utilize dense feature representations, improving model performance in both reasoning and explaining. We propose a novel model named Feature-Enhanced Neural Collaborative Reasoning (FENCR) in an end-to-end manner. Firstly, we construct dense feature embeddings with a selection module to automatically generate distinct user and item feature groups. Then, neural logical modules are designed to conduct logical operations (i.e., conjunction and disjunction) over the selected feature groups, by which logical rules are generated. Self-supervised logical axiom regularizers are accordingly introduced in the loss function to augment the logical properties of the modules. Finally, recommendation results are derived by personalized aggregation of multiple generated logical rules. Comparative experiments on two large public datasets show our model outperforms state-of-the-art logic-enhanced recommendation models. Further analyses also showcase the explainability of the model facilitated by derived rules.
Index Terms:
collaborative reasoning, interpretable modeling, explainable recommendation, rule learning
I Introduction
Explainability has always been a topic of great concern within the field of recommendation [ExplainSurveyYongfengZhang, chen2022measuring, sun2020dual, sun2021unsupervised]. Researchers explore various methods to help users understand why recommendation systems give certain results. Among these methods, the text-based explanation emerges as a kind of prominent and widely-used approach [DBLP:GenerateLanguage, li2021personalizedTransformer, li2022personalized]. Through explanation text, systems can present the reason for recommendations to users in natural language, thereby increasing user trust and comprehensibility of the given results.
An effective evaluation should ensure the explanations truly resonate with users and meet their expectations. However, with advanced approaches developed for generating explanatory text, assessing their quality remains an issue that has yet to be adequately resolved. Existing methods for evaluation can be categorized into three main types: self-report, third-party annotations, and reference-based metrics. Conducting user studies to obtain self-reported feedback can most accurately reflect user experience. This approach requires evaluations to be recorded alongside the recommendations and is, therefore, difficult to obtain and use in public datasets. Third-party annotations can reflect human feedback and are relatively accessible. Still, manual labeling is expensive, time-consuming, and lack of scalability. Reference-based metrics assess the quality of the target text based on the reference text and offer a standard assessment that is relatively easy to acquire. Common metrics such as BLEU [DBLP:bleu2022ACL], and ROUGE [Rouge2004Lin] calculate the similarity between the generated text and the reference text. However, there may not exist reference texts for the scenario (many use reviews as a substitute). In addition, these textual similarities metrics may not ideally reflect user perception of recommendation explanation [ExpScore2022WWW, stopbleu]. These limitations highlight the need to develop evaluation methodologies that are in-line with human experience, easy to acquire, and reproducible.
Recently, the development of large language models (LLM) sheds new light on the evaluation of various neural language generation (NLG) tasks [kocmi2023LLMEvaluateTransaltion, wang2023automatedpersonaliztion, wang2023chatgpt, fu2023gptscore]. Considering that LLMs can follow human instructions and their language modeling ability, LLMs can make adequate evaluations under reference-free settings with appropriate prompt [GEval2023liu, he2024annollm]. LLM offers an appealing solution for evaluating the quality of recommendation explanations since it is efficient (lower cost than manual labeling) and widely applicable (almost no dataset limitation). In a study on the design of explainable recommendation method [lei2023recexplainer], researchers introduce LLMs to evaluate the quality of explanations generated. However, we state that successes in the evaluation of general NLG tasks can not be migrated to the evaluation of explainable recommendations without verification. Unlike those NLG tasks that are already been examined, evaluating recommendation explanation text is more sophisticated. The reason is that measuring the quality of recommendations involves a group of diverse goals, e.g. persuasiveness, transparency [ConflictGoalsOfExplanations], etc. Additionally, a large portion of these goals are related to the subjective perception of users. All these factors add to the difficulty of assessing the quality of recommendation explanation texts. These considerations underscore the need to further explore the feasibility of leveraging LLM as a potential solution for text-based explainable recommendation evaluation.
In this paper, we are concerned with the research question: Can LLM serve as an evaluator of recommendation explanation text? In particular, we delve into three detailed inquiries:
RQ1 Can LLMs evaluate different aspects of user perceptions about the recommendation explanation texts in a zero-shot setting?
RQ2 Can LLMs collaborate with human labels to enhance the effectiveness of evaluation?
RQ3 Can LLMs collaborate with each other to enhance the effectiveness of evaluation?
To answer these research questions, we use data from a user study in previous work [UserPerceptionTois], including the users’ ratings on 4 aspects of the provided explanatory texts, as well as the self-explanation text of the users for the recommended movies. Our study is based on the premise that the user receiving the recommendation is the ground-truth evaluator for the quality of the explanation. To study the reliability of LLM annotations, we additionally collect third-party annotations and LLM annotations under human instructions. To comprehensively compare between evaluators, we design and apply a 3-level meta-evaluation strategy to measure the correlation between evaluator annotations and user labels on different aspects. We compare the evaluation accuracy of LLM evaluators with third-party annotations and commonly used reference-based metrics, i.e., BLEU [DBLP:bleu2022ACL] and ROUGE [Rouge2004Lin].
Our main findings are: (1) Certain zero-shot LLMs, such as GPT4, can attain evaluation accuracy comparable to or better than traditional methods, with performance varying across different aspects. (2) The effectiveness of one-shot learning depends on backbone LLMs. Particularly, personalized cases can assist GPT4 in learning user scoring bias. (3) Ensembling the scoring of multiple heterogeneous LLMs can improve the accuracy and stability of evaluation.
In summary, our contributions include investigating the feasibility of using zero-LLM as the evaluator for explainable recommendations, discussing possible collaboration paradigms between LLM and human labels, and exploring the aggregations of multiple LLM evaluators. We propose that LLM can be a reproducible and cost-effective solution for evaluating recommendation explanation text with appropriate prompts and settings. Compared with traditional methods, LLM-based evaluators can be applied to new datasets with few limitations. By introducing this evaluation approach, we aspire to contribute to the advancement of the area of explainable recommendation.
II Experimental Setup
Figure 1: Traditional evaluation approaches vs. utilizing LLMs for evaluations.
II-A Problem Formulation
We use $\mathcal{U}=(u_{1},u_{2},...u_{|\mathcal{U}|})$ to denote the set of users in an RS (Recommendation System). The RS recommend multiple items to each user $u$, which are defined as $\mathcal{I}_{u}=(i_{1},i_{2},...,i_{|\mathcal{I}_{u}|})$. When the item $i\in\mathcal{I}_{u}$ is recommended to the user $u$, explanation texts are generated by a group of generation methods, which are denoted as $\mathcal{G}=(g_{1}(\cdot),g_{2}(\cdot),...,g_{|\mathcal{G}|}(\cdot))$. $E_{u,i}=g(u,i)$ denotes the explanation text given by $g$ when recommending item $i$ to user $u$.
$f(\cdot)$ denotes the evaluation methods for $E$. We assume that users in the system are the most accurate evaluators for explanations of items recommended to them. Their evaluations are denoted as $\mathbf{s_{(u,i,E)}}=f_{u}(i,E)$. When utilizing $f_{LLM}$ to approximate $f_{u}$, the evaluation given by LLM can be represented as:
$\hat{\mathbf{s}}_{LLM(u,i,E)}=f_{LLM}(u,i,E)=LLM(u,i,g(\cdot),\mathcal{P})$
where $\mathcal{P}$ denotes the prompt that contains human instructions.
Meta-evaluation method $h(\cdot)$ based on correlation metrics are introduced to measure the similarity between $\mathbf{s}$ and $\hat{\mathbf{s}}$. The accuracy of the evaluation given by LLM can be expressed as $h(\mathbf{s},\hat{\mathbf{s}}_{LLM})$. In our paper, we examine the feasibility of $f_{LLM}$ by analysing the value of $h(\mathbf{s},\hat{\mathbf{s}}_{LLM})$. $h(\mathbf{s},\hat{\mathbf{s}}_{t})$ and $h(\mathbf{s},\hat{\mathbf{s}}_{r})$ are used as referent standards, where $\hat{\mathbf{s}}_{t}$ denotes evaluation given by third-party annotators, and $\hat{\mathbf{s}}_{r}$ denotes evaluation given by referenece-based metrics.
II-B Data Construction
II-B1 Data overview
[…]
UserPerceptionTois create a movie recommendation platform. It first captures user preferences, and then gives personalized recommendations along with text-formed explanations generated by a series of systems. We utilize two parts of the data: 1) self-explanations written by users, 2) 1-5 ratings of users to explanations generated by different methods in terms of 4 aspects, which are persuasiveness, transparency, accuracy, and satisfaction.
Formally, 39 participants are included in $\mathcal{U}$ and $\mathcal{I}$ are movies from Movielens Latest dataset [^1] https://grouplens.org/datasets/movielens/latest/. $\mathcal{I}_{u}$ denotes the itemset recommended to the user $u$, which includes the top-8 movies calculated by BiasedMF [RendleMF2012]. $\mathcal{G}$ includes a series of systems used to generate the explanatory texts. Most of these methods are template-based methods, e.g. user-based method, which generates explanations in the form of: “[N%] of users who share similar watching tastes with you like [MOVIE TITLE] after watching it.” $\mathcal{G}$ also includes the system that directly generates the complete natural language sentence, i.e., peer-explanation written by others. The data includes the self-explanation for each user-item pair. Since they are written by users themselves, there is no evaluation rating attached.
In summary, the data comprises entries from 39 users, each of whom received recommendations for 8 movies. For each user-movie pair, approximately 8 explanations are generated, resulting in a total of around 2,500 text entries. The data are collected in Chinese. We take this data as ground truth $\mathbf{s}$ in the experiments. We additionally collect evaluation scores $\mathbf{\hat{s}}$ given by third-party annotators, LLMs and quantitative metrics. The evaluation approaches from which the data is derived are illustrated in Figure 1.
II-B2 Evaluations from users
For each explanation text $E$ generated, the user is asked to give a 5-scale Likert score to the explanations. The user feedback of the explanation is given on 4 aspects:
Persuasiveness: This explanation is convincing to me;
Transparency: Based on this explanation, I understand why this movie is recommended;
Accuracy: This explanation is consistent with my interests;
Satisfaction: I am satisfied with this explanation.
These 5-scale Likert questions are also utilized in collecting third-party evaluations and LLM evaluations to ensure consistency in the definitions of aspects. The user annotation for each explanatory text $\mathbf{s}_{(u,i,E)}$ is a one-dimensional vector of length 4, with elements being integers between 1 to 5.
II-B3 Evaluation from third-party annotators
We employ two annotators to evaluate the explanatory text from the above-mentioned four aspects. The annotators are informed that the user is utilizing a movie recommendation platform and receiving film recommendations along with the reasons for those recommendations. The annotators are asked to score the explanatory text in an item-wise manner. The third-party annotation for each explanatory text $\mathbf{\hat{s}}_{t(u,i,E)}$ is a one-dimensional vector of length 4, with elements being integers between 1 to 5.
II-B4 Evaluation from Quantitative metrics
We utilize BLEU [DBLP:bleu2022ACL] and ROUGE [Rouge2004Lin] to evaluate the explanatory text, taking the self-explanations written by users as reference texts. The evaluation score $\mathbf{\hat{s}}_{r}$ is a float that measures text similarity.
II-C LLM as Evaluator
II-C1 LLM Evaluator Construction
We utilize pre-trained LLMs to provide annotations for each explanation text $E$. The large language model receives the movie name and the corresponding explanation text, accompanied by a prompt $\mathcal{P}$ to describe the context. Although $E$ is in Chinese as introduced in Section II-B1, we apply $\mathcal{P}$ in English to ensure a standardized prompt design. The LLM evaluation for each explanatory text $\mathbf{\hat{s}}_{LLM(u,i,E)}$ is a one-dimensional vector of length 4, with elements being integers between 1 to 5. Our experiments are based on item-wise evaluation, where only one text entry is evaluated at a time. This approach aligns with the setup in the user study.
To investigate the aforementioned RQs, three basic methods are used to construct the LLM evaluator, as illustrated in Figure 1. First, we directly use zero-shot LLMs for evaluation. Then, we consider providing human labels as contextual information to enhance their abilities to learn user subjective perceptions. Finally, inspired by the traditional approach of collecting multiple annotators labels and then averaging the results, we similarly ensemble the results of multiple annotations. When constructing LLM evaluators, we aim to maintain the simplicity and transferability of the approaches. Therefore, we have not leveraged other common methods that require model training, such as fine-tuning.
Figure 2: The outline of evaluation prompt templates applied in our study.
II-C2 Prompt Construction
In the section, we mainly introduce how we construct the prompt $\mathcal{P}$. The prompt is designed to guide the LLM in evaluating the quality of explanatory text from specified aspects. It includes several key components, briefed in Figure 2. We attempted to integrate user information in the prompt but found that this approach did not improve performance and sometimes even decreased it. Following, we describe in detail the different settings in prompt constructions.
Single-Aspect vs. Multiple-Aspect. The difference between single and multiple-aspect prompts lies in the Aspect part. A multiple-aspect prompt instructs the model to evaluate all four aspects concurrently, while a single-aspect prompt focuses on one aspect at a time. Utilizing single-aspect prompts necessitates more times of interactions with LLM compared to multiple-aspect prompts. Experimental results under the two settings are presented and compared in RQ1, Section II-D.
{algorithm}
Comparison of personalized and non-personalized one-shot learning process
PersonalizedOneShot each user $u$in $\mathcal{U}$Randomly choose item $i_{0}$from $\mathcal{I}_{u}$ollect user evaluations on explanatory text $E=g(u,i_{0})$as $\mathbf{s_{0}}_{g}$for each $g$in $\mathcal{G}$each movie $i$in $\mathcal{I}_{u}$, each $g$in $\mathcal{G}$$Prompt\leftarrow$ConstructPrompt($i$, $g(u,i)$, $\mathbf{s_{0}}_{g}$) $Result\leftarrow$LLM($Prompt$) Parse $Result$and process the answer Non-PersonalizedOneShot Randomly choose user-item pair $(u_{0},i_{0})$Collect user evaluations on explanatory text $E=g(u_{0},i_{0})$as $\mathbf{s_{0}}_{g}$for each $g$in $\mathcal{G}$each user $u$in $\mathcal{U}$,each movie $i$in $\mathcal{I}_{u}$, each $g$in $\mathcal{G}$$Prompt\leftarrow$ConstructPrompt($i$, $g(u,i)$, $\mathbf{s_{0}}_{g}$) $Result\leftarrow$LLM($Prompt$) Parse $Result$and process the answer
[…]
Zero-Shot vs. One-Shot. Zero-shot learning presents the task without any example and relies solely on the pre-trained knowledge and reasoning ability of the LLM. While this may result in a lack of context-specific guidance, the advantage is that it can be applied directly to datasets without manual labeling.
To explore whether humans and LLM can collaborate on evaluation, we investigate the impact of utilizing human labels as contextual information in LLM evaluation. Our primary focus is on one-shot learning and includes one human label in the prompt. To enhance the LLM’s ability to learn personalized preferences, we employ personalized one-shot learning. This approach involves providing the scoring example from the same user for the target data. Formally, for the user $u$, item $i$ and explanation $E=g(u,i)$, the personalized example is the scoring given to $E^{\prime}=g(u,i_{0})$, which is the explanation generated by the same system $g$ and the same user $u$ on a randomly chosen movie item $i_{0}$. The methods help the prompt $\mathcal{P}$ incorporate personal information. However, a limitation is that it still requires collecting real user feedback, which is expansive and sometimes impractical. Hence, we also investigate non-personalized one-shot prompts. In this setting, all examples are from randomly selected $(u_{0},i_{0})$ pair. That is, the example can come from another user and are easier to collect in practice. Formally, for the user $u$, item $i$ and explanation $E=g(u,i)$, the non-personalized example is the scoring given to $E^{\prime}=g(u_{0},i_{0})$. We summarize the personalized and non-personalized one-shot learning procedures in Algorithm 1 to better illustrate the example selection process and their differences. Experimental results under different example settings are presented and compared in RQ2, Section LABEL:sec:RQ2.
II-D Three-Level Meta Evaluation
Good evaluation requires that evaluation efforts themselves be evaluated [stufflebeam1974metaevaluation]. Employing the suitable meta-evaluation method $h(\cdot)$ is crucial for thoroughly examining the evaluation procedure. Correlation coefficients, such as Pearson($r$), Spearman($\rho$), and Kendall($\tau$) coefficients, have served as a widely used metric for gauging the similarity in trends between two arrays of ratings or scores [liu2023calibrating]. In NLG, the strategy of meta-evaluation can be divided into two levels: Dataset-Level and Sample-Level [liu2023calibrating]. Formally, given a dataset $\mathcal{D}$ consisting of a set of source data $\mathcal{X}=(x_{1},x_{2},....,x_{|\mathcal{X}|})$ and generation methods $\mathcal{G}=(g_{1}(\cdot),g_{2}(\cdot),...,g_{|\mathcal{G}|}(\cdot))$ and correlation metric $r(\cdot)$ dataset-level and sample-level meta evaluation are expressed as:
Dataset Level
$h_{D}(\mathbf{s},\hat{\mathbf{s}})=r\left(\left(\mathbf{s}_{1},\mathbf{s}_{2},...\mathbf{s}_{|\mathcal{X}||\mathcal{G}|}\right),\left(\hat{\mathbf{s}}_{1},\hat{\mathbf{s}}_{2},...\hat{\mathbf{s}}_{|\mathcal{X}||\mathcal{G}|}\right)\right)\$
Sample Level
$h_{S}(\mathbf{s},\hat{\mathbf{s}})=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}r\left(\left(\mathbf{s}_{(i,1)},...\mathbf{s}_{(i,|\mathcal{G}|)}\right),\left(\hat{\mathbf{s}}_{(i,1)},...\hat{\mathbf{s}}_{(i,|\mathcal{G}|)}\right)\right)$
Previous studies have not discussed a multiple-level meta-evaluation when assessing the quality of evaluation metrics for explainable recommendations. However, considering multiple-level meta-evaluation in explainable recommendations is worthwhile. This is because the distribution of ground-truth labels or evaluations generated by models may differ between users or user-item pairs. Consequently, while an evaluation metric might effectively capture trends within specific groups (such as comparing the qualities of a group of texts derived from the same user-movie pair), it may struggle to accurately depict trends across groups (such as capturing certain users’ inclination to provide higher ratings).
Therefore, we propose a 3-level strategy for meta-evaluation on recommendation explanations. Our motivation is to divide the data into groups at various grain scales and measure the correlation between evaluation results with ground-truth labels within each group. The three proposed levels are Dataset-Level, User-Level, and Pair-Level. Together, they provide a comprehensive view for $h$:
Dataset Level correlation calculates the correlation of all scores. The total number of data can be expressed as $|\mathcal{D}|=|\mathcal{U}|\cdot|\mathcal{I}_{u}|\cdot|\mathcal{G}|$.
$h_{D}(\mathbf{s},\hat{\mathbf{s}})=r\left(\left(\mathbf{s}_{1},\mathbf{s}_{2},...\mathbf{s}_{|\mathcal{D}|}\right),\left(\hat{\mathbf{s}}_{1},\hat{\mathbf{s}}_{2},...\hat{\mathbf{s}}_{|\mathcal{D}|}\right)\right)$
User Level correlation calculates the correlation within the data of each user and then averages it. The total number of data derived from the same user $u$ can be expressed as $|\mathcal{D}_{u}|=\cdot|\mathcal{I}_{u}|\cdot|\mathcal{G}|$.
$h_{U}(\mathbf{s},\hat{\mathbf{s}})=\frac{1}{|\mathcal{U}|}\sum_{u}r\left(\left(\mathbf{s}_{1},\mathbf{s}_{2},...\mathbf{s}_{|\mathcal{D}_{u}|}\right),\left(\hat{\mathbf{s}}_{1},\hat{\mathbf{s}}_{2},...\hat{\mathbf{s}}_{|\mathcal{D}_{u}|}\right)\right)$
Pair Level correlation calculates the correlation within the data of each user-item pair and then averages it. $|\mathcal{G}|$ explanation texts are derived from the same user-item pair by $g_{1}(\cdot),g_{2}(\cdot),...,g_{\mathcal{G}}(\cdot)$.
$h_{P}(\mathbf{s},\hat{\mathbf{s}})=\frac{1}{|\mathcal{U}|\cdot|\mathcal{I}_{u}|}\sum_{u,i\in\mathcal{I}_{u}}r\left(\left(\mathbf{s}_{1},\mathbf{s}_{2},...\mathbf{s}_{|\mathcal{G}|}\right),\left(\hat{\mathbf{s}}_{1},\hat{\mathbf{s}}_{2},...\hat{\mathbf{s}}_{|\mathcal{G}|}\right)\right)$
Referring to correlations on which level depends on the settings of the task. For instance, if an evaluation metric is used to compare between $\mathcal{G}$, it is more important to look at the Pair-Level correlation between the results from the evaluation metric and ground-truth labels. When the study requires a measure of satisfaction with the recommendation explanation text by different users of the system, the results of Dataset-Level correlation should be referred to.
We use a concrete example to illustrate the possible variation of the same evaluation metric at different levels. For instance, our experimental results in Section II-D found that BLEU-4 scoring results correlate poorly or even negatively with ground-truth labels at the Dataset-Level, whereas correlations improve at User-Level and Pair-Level. This discrepancy arises because BLEU-4 computes the token similarity between the target text and the reference text. Some reference texts may contain more commonly used token combinations than others, potentially inflating the BLEU-4 score. This introduces a bias into the evaluation process that is unrelated to user perceptions, thereby impacting the Dataset-Level correlation. Full experimental results and analyses can be found in Section II-D.
TABLE I: The 3-level Pearson correlation between the results given by the evaluator and the ground-truth label given by users. Bold fonts denote the best results among all tested evaluators and the underlines show the second-best results.
Dataset-Level / User-Level / Pair-Level (%)
Method Persuasiveness Transparency Accuracy Satisfaction Average
Random -0.55 / 0.52 / 1.81 0.65 / -0.43 / -2.58 -0.41 / 4.12 / 3.98 0.36 / -2.26 / 5.88 0.01 / 0.49 / 2.27
Reference-based Metric
BLEU-1 11.68 / 15.84 / 17.07 10.06 / 12.69 / 14.44 6.43 / 10.71 / 12.18 11.36 / 12.91 / 15.79 9.88 / 13.04 / 14.87
BLEU-4 -1.17 / 7.68 / 13.53 -3.47 / 4.13 / 10.24 -4.63 / 4.8 / 8.96 0.61 / 6.86 / 12.09 -2.16 / 5.86 / 11.21
ROUGE-1-F 14.16 / 16.39 / 17.56 11.93 / 12.74 / 14.45 8.61 / 11.02 / 12.83 12.87 / 13.2 / 16.23 11.89 / 13.34 / 15.27
ROUGE-L-F 14.16 / 16.39 / 17.56 11.93 / 12.74 / 14.45 8.61 / 11.02 / 12.83 12.87 / 13.2 / 16.23 11.89 / 13.34 / 15.27
Annnotation
Annotator-1 19.88 / 18.31 / 16.72 15.66 / 16.18 / 11.31 10.16 / 9.78 / 9.77 14.93 / 13.28 / 12.69 15.16 / 14.39 / 12.62
Annotator-2 21.4 / 21.17 / 20.9 25.97 / 26.42 / 27.84 10.96 / 10.96 / 9.32 8.86 / 9.72 / 9.43 16.8 / 17.07 / 16.87
Average 23.33 / 22.25 / 20.93 24.53 / 25.36 / 23.12 12.83 / 12.52 / 11.19 13.9 / 13.54 / 13.16 18.65 / 18.42 / 17.10
Single-Aspect Prompt
Llama2-7B -4.02 / -3.32 / -3.9 -1.52 / -2.92 / -5.43 -1.11 / -1.88 / -3.76 0.74 / 2.72 / 4.54 -1.48 / -1.35 / -2.14
Llama2-13B 8.39 / 9.5 / 10.91 10.64 / 11.67 / 10.68 -4.44 / -4.52 / -0.96 -0.18 / 1.12 / 0.94 3.60 / 4.44 / 5.39
Qwen1.5-7B 5.81 / 8.14 / 10.78 5.49 / 5.07 / 6.15 6.26 / 6.35 / 5.07 -1.97 / -1.52 / -2.42 3.9 / 4.51 / 4.89
Qwen1.5-14B 7.13 / 6.92 / 7.01 22.61 / 22.75 / 22.33 28.65 / 30.71 / 35.11 13.88 / 13.68 / 13.94 18.07 / 18.52 / 19.60
GPT3.5-Turbo 26.81 / 26.36 / 29.58 20.62 / 21.22 / 25.01 16.33 / 15.56 / 17.93 9.95 / 7.75 / 6.33 18.43 / 17.72 / 19.71
GPT4 18.36 / 19.78 / 22.03 20.17 / 21.57 / 23.62 14.46 / 15.61 / 14.33 7.49 / 5.92 / 3.17 15.12 / 15.72 / 15.79
Multiple-Aspect Prompt
Llama2-7B -1.26 / -2.85 / -14.34 -2.2 / -2.59 / -8.87 -3.36 / -7.23 / -16.36 1.74 / 1.99 / 1.82 -1.27 / -2.67 / -9.44
Llama2-13B 17.04 / 17.33 / 18.56 4.26 / 3.41 / 10.25 3.59 / 2.1 / 2.24 17.93 / 16.82 / 18.52 10.71 / 9.92 / 12.39
Qwen1.5-7B 13.0 / 13.26 / 13.08 11.75 / 11.74 / 15.28 -0.8 / -0.34 / -0.49 10.63 / 9.28 / 15.6 8.65 / 8.49 / 10.87
Qwen1.5-14B 25.85 / 26.53 / 32.28 18.16 / 18.45 / 22.03 12.25 / 11.32 / 15.26 15.82 / 14.83 / 18.26 18.02 / 17.78 / 21.96
GPT3.5-Turbo 26.41 / 26.36 / 28.2 11.16 / 9.86 / 11.38 12.09 / 10.63 / 11.15 20.93 / 19.56 / 20.78 17.65 / 16.61 / 17.88
GPT4 27.26 / 28.25 / 28.99 12.68 / 12.17 / 13.26 20.30 / 22.04 / 24.93 24.05 / 25.12 / 27.35 21.07 / 21.90 / 23.63
(a) GPT4 (M)
(b) Qwen1.5-14B (M)
Figure 3: Comparison of 3-level Pearson correlations for zero-shot, (non-personalized) one-shot, and personalized one-shot learning on GPT4(M) and Qwen(M). (M) denotes the multiple-aspect version.
III Conclusion
In this paper, we propose FENCR, which learns logical rules as explanations and feature embedding for representation. We design novel neural logical modules that represent conjunction and disjunction to better learn feature interactions in logic operations on vectors. The inference procedure of FENCR  can be interpreted into logical rules. Experiments show FENCR  outperforms previous logic-enhanced recommendation models while being able to provide rule-based reasons for the results. We believe that using logic rules as explanations is a topic worth exploring. As explanations in the form of logic rules are easy to understand and modify, they are more stable and controllable when supporting downstream applications. In the future, we will further study the potential downstream tasks of generated rules, including providing guidance to feature engineering and collaborating with language models for generating more persuasive texts.